If the concentration of a solution of NaOH is 4g/dm3 ,then what mass of NaOH is contained in 250 cm3 of this solution ? [1 dm3 = 1000 cm3 ]?
1
Expert's answer
2014-05-08T03:46:33-0400
Express concentration in g/cm3:
c = 4g/dm3 = 4g/1000 cm3 = 0.004 g/cm3
Find the mass of NaOH:
m = c*V = 250 cm3 * 0.004 g/cm3 = 1g